# Efficient Preimage Approximation for Neural Network Certification

### v1

@@ Abstract
The growing reliance on artificial intelligence in safety- and security-critical applications demands effective neural network certification. A challenging real-world use case is certification against “patch attacks”, where adversarial patches or lighting conditions obscure parts of images, for example traffic signs. One approach to certification, which also gives quantitative coverage estimates, utilizes preimages of neural networks, i.e., the set of inputs that lead to a specified output. However, these preimage approximation methods, including the state-of-the-art PREMAP algorithm, struggle with scalability. This paper presents novel algorithmic improvements to PREMAP involving tighter bounds, adaptive Monte Carlo sampling, and improved branching heuristics. We demonstrate efficiency improvements of at least an order of magnitude on reinforcement learning control benchmarks, and show that our method scales to convolutional neural networks that were previously infeasible. Our results demonstrate the potential of preimage approximation methodology for reliability and robustness certification.
[…]
3853
@@ 1 Introduction
The widespread deployment of AI technologies places significant demands on their trustworthiness and reliability, which are of particular concern in high-stakes applications, as recognized in the EU AI Act [9]. Much effort has been devoted to ensuring safety and security of AI applications involving neural networks (NNs) because of their instability to adversarial perturbations [12]. A particularly challenging real-world example are “patch attacks” on images, such as traffic signs. In Figure 1, we show two examples of physical patch attacks: real graffiti on a traffic sign [10] (left) and a patch caused by a sunbeam [5] (middle), which can result from a specific adversarial manipulation of the image or for natural reasons.
@@
The above examples highlight the importance of studying robustness of image recognition models to adversarial patch attacks for applications such as safe autonomous driving or biometric security software. To this end, a wide range of methodologies have been developed for neural networks to enhance their robustness, either via (certified) adversarial training [26, 20] or designing certifiable [33] or empirical [31] defences. An alternative method is to certify robustness by computing deterministic [13, 15] or probabilistic [30, 36] guarantees on the network outputs. A drawback of certifiable defences is their relatively high inference time, although there are viable trade-offs [33, 32], Similarly to formal verification, certification is more powerful than adversarial training or empirical defences because it can guarantee that the network output is correct for all inputs in a given region. Using the patch attack in Figure 1 (right) as an illustration, and assuming the patch is placed in a fixed position, certification aims to guarantee the prediction for all possible (RGB) colour values.
@@
Adversarial robustness certification methods commonly work by bounding the outputs of the neural network for a given input subdomain, which may yield loose bounds for the resulting output over-approximation. A recently developed alternative approach focuses instead on bounding the preimage of a given output specification (typically a polyhedron), either through over-approximating the preimage [16, 39] or under-approximating it [38, 39]. Preimage under-approximation, in contrast to over-approximation, offers stronger compliance guarantees, since all inputs in the computed approximation satisfy the guarantee, and quantitative verification, i.e., estimating the proportion of inputs that meet the specification, which can directly serve as a reliability certificate. However, despite the positive experimental results reported in [38, 39], existing methods do not scale to high dimensions, such as commonly used convolutional neural networks, which is compounded by the prohibitive size of the branching computation tree.
In this paper, we aim to significantly improve the scalability and computational performance of the state-of-the-art preimage approximation algorithm PREMAP [38, 39]. This raises a number of challenges, which we resolve by careful analysis that allowed us to select effective algorithmic improvements such as bound tightening, optimized branching and refinement, as well as adaptive Monte Carlo sampling. We experimentally evaluate our method on a range of case studies, showing that our method is able to complete almost twice the number of cases solved using the original method. Now we are able to certify robustness to patch attacks for convolutional networks applied on GTSRB (traffic signs), SVHN (house numbers) and CIFAR (standard image benchmark), which were not feasible with the original PREMAP. We also provide a comparison with the previous version of PREMAP on controller benchmarks, showing at least an order of magnitude improvements in running time. Additionally, we demonstrate how PREMAP can be used for interpretability, to investigate important regions of the input. Our results demonstrate the potential of preimage approximation methodology for exploitation in reliability and robustness certification.
2 Background
@@
This section presents the relevant background. In Section 2.1, we define preliminary concepts and provide an overview of the PREiMage APproximation (PREMAP) method [38, 39], focusing on key aspects needed for our algorithmic improvements; for more detail of the techniques please refer to [39]. Related works are discussed in Section 2.2.
2.1 Preliminaries
Neural networks.
@@ 2.1 Preliminaries
We use $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ to denote a feed-forward neural network, where $f(\mathbf{x})=\mathbf{y}$. For layer $l$, $\mathbf{W}^{l}$ denotes the weight matrix, $\mathbf{b}^{l}$ the bias, $\mathbf{z}^{l}$ the pre-activation neurons, and $h^{l}(\mathbf{z}^{l})$ the activation function, such that $\mathbf{z}^{l}=\mathbf{W}^{l}h^{l-1}(\mathbf{z}^{l-1})+\mathbf{b}^{l}$. We use under- and overlines to denote the lower and upper bounds, e.g., $\underline{\mathbf{x}}_{i}\leq\mathbf{x}_{i}\leq\overline{\mathbf{x}}_{i}$. In this paper we focus on ReLU neural networks, where $h^{l}(\mathbf{z}^{l})=\text{ReLU}(\mathbf{z}^{l})=\max(\mathbf{z}^{l},0)$ is applied element-wise, but our method can also be generalized to other activation functions that can be bounded by linear functions [37, 6].
Linear relaxation.
@@
For the neural network $f$, linear relaxation is used to compute linear lower and upper bounds for the output like
@@
The anytime PREMAP algorithm [39, 38], implements both under- or over-approximation of the preimage of a given output specification $\mathbb{O}$ restricted to the input domain $\mathbb{I}$, in the form of a disjoint union of polytopes. However, in this paper we will focus on under-approximations because they offer stronger guarantees and enable quantitative verification. Algorithm 1 shows an overview of the method with our modification highlighted with green shading (discussed further in Section 3). PREMAP relies on linear relaxation with LiRPA to approximate the preimage (see 3), which is further refined through splitting the problem into disjoint subregions (on 8 and 9), which can be approximated independently in parallel (see 10 and 11) with a divide-and-conquer approach. The algorithm terminates upon reaching a preset threshold for the preimage approximation, with exact rather than statistical quantitative guarantees.
Branching and refinement.
@@
Since the initial bound given by LiRPA might be very loose (Algorithm 1, 3), PREMAP iteratively partitions a subdomain $\mathbb{I}_{\text{sub}}\subseteq\mathbb{I}$ into two smaller subdomains $\mathbb{I}_{\text{sub}}^{-}$ and $\mathbb{I}_{\text{sub}}^{+}$ (5 to 12). Which subdomain to split is selected by the largest difference in estimated volume between the exact preimage and the approximation in the subdomain (lines 6). The split can be either on the input, which partitions the input into two subdomains along some feature, or based on the pre-activation value of an intermediate unstable neuron [39], called ReLU splitting. As this paper deals with larger input domains, for which input splitting performs less well, we will focus on the latter (8).
@@
A key advantage of ReLU splitting is that it allows us to replace unstable neuron bounds with precise bounds. For an unstable ReLU neuron $h^{l}(\mathbf{z}^{l}_{j})=\max(0,\mathbf{z}^{l}_{j})$, we use linear relaxation to bound the post-activation value as in Equation 1. When a split is introduced, the neuron becomes stable in each subdomain, and the exact linear functions $h^{l}(\mathbf{z}^{l}_{j})=\mathbf{z}^{l}_{j}$ and $h^{l}(\mathbf{z}^{l}_{j})=0$ can be used in place of its linear relaxation. This can typically tighten the approximation on each subdomain as the linear relaxation errors for this unstable neuron are removed and substituted with the exact symbolic function.
@@
The bounds, $\underline{\mathbf{A}}\mathbf{x}+\underline{\mathbf{b}}$, for the new under-approximations are optimized via projected gradient descent (Algorithm 2, 3) and the process of refining the subdomains continues until an appropriate termination criterion is met (Algorithm 1, 7). The procedure can also be stopped anytime to produce a valid (but smaller) approximation. The (anytime) union of the preimage approximations $\mathbf{P}$ for the disjoint subdomains $\mathbb{I}_{\text{sub}}$ constitutes the preimage approximation of the output property.
Algorithm 1 Premap is an algorithm for finding preimage under approximations. The inputs are the neural network $f$, the input domain $\mathbb{I}$, the output specifications $\mathbb{O}$, the early stopping threshold $v_{\text{th}}$, and the number of samples $n$. Some functions are detailed in Algorithm 2. The lines with our modifications compared to [39] are highlighted with green shading and are discussed in detail in Sections 3.2, 3.1 and 3.3.
@@
2:     $\mathbf{X}\leftarrow\textsc{Sample}(\mathbb{I},5n)$$\rangle$ ⟩ Sample extra when it is easy
@@
4:     $\mathbb{D}\leftarrow\{(\mathbf{P},\mathbb{I},\mathbf{X})\}$$\rangle$ ⟩ Approximation, domain, samples
[…]
@@
7:         $\mathbf{X}\leftarrow\textsc{Sample}(\mathbb{I}_{\text{sub}},n,\mathbf{X})$$\rangle$ ⟩ Top up samples
8:         $l,i\leftarrow\textsc{SelectNeuron}(f_{\mathbb{O}},\mathbb{I}_{\text{sub}},\mathbf{X})$
9:         $\mathbb{I}_{\text{sub}}^{-},\mathbf{X}^{-},\mathbb{I}_{\text{sub}}^{+},\mathbf{X}^{+}\leftarrow\textsc{SplitNeuron}(\mathbb{I}_{\text{sub}},\mathbf{X},l,i)$
10:         $\mathbf{P}^{-}\leftarrow\textsc{PreimageApprox}(f_{\mathbb{O}},\mathbb{I}_{\text{sub}}^{-},\mathbf{X}^{-})$
11:         $\mathbf{P}^{+}\leftarrow\textsc{PreimageApprox}(f_{\mathbb{O}},\mathbb{I}_{\text{sub}}^{+},\mathbf{X}^{+})$
12:         $\mathbb{D}\leftarrow\mathbb{D}\cup\{(\mathbf{P}^{-},\mathbb{I}_{\text{sub}}^{-},\mathbf{X}^{-}),\leavevmode\nobreak\ (\mathbf{P}^{+},\mathbb{I}_{\text{sub}}^{+},\mathbf{X}^{+})\}$
13:     return $\{\mathbf{P}\mid(\mathbf{P},\cdot,\cdot)\in\mathbb{D}\}$
Algorithm 2 Additional functions for Premap, see Algorithm 1 for details. The LIRPAfunction produces linear bounds using the LiRPA library [34] with optimisable parameters $\boldsymbol{\alpha}$.
@@
3:     $\boldsymbol{\alpha},\beta\leftarrow\operatorname*{arg\,max}\limits_{\boldsymbol{\alpha}}\sum\limits_{\mathbf{x}\in\mathbf{X}}\sigma(-\log(\sum\exp({-\underline{\mathbf{A}}^{(\boldsymbol{\alpha})}\mathbf{x}-\underline{\mathbf{b}}^{(\boldsymbol{\alpha})}})))$
@@
5:     return $\mathbf{P}$$\rangle$ ⟩ Preimage under approximation
6:function EstimateCoverage($\mathbb{D},f_{\mathbb{O}}$)
@@
9:     return $v_{\mathbf{P}}/v_{\mathbb{O}}$$\rangle$ ⟩ Approximated fraction of the preimage
@@
12:     return $v_{\mathbb{O}\setminus\mathbf{P}}$$\rangle$ ⟩ Volume of preimage not approximated
2.2 Related works
Robustness certification methods focus on computing deterministic [13, 15] or probabilistic/statistical [30, 36] guarantees on the output of a neural network. The methods can be classified into complete methods, such as constraint solving [15], or sound though incomplete methods, e.g., convex relaxation [25], which can be strengthened to ensure completeness by employing branch-and-bound procedures [3, 4]. Building on [38, 39], our work employs convex (linear) relaxation and Monte Carlo sampling in a divide-and-conquer framework for approximating the preimage of the neural network with exact (not statistical) quantitative guarantees. We leverage LiRPA [34] as implemented in CROWN tools [37, 35, 29] to compute symbolic relaxation bounds, adapting to preimage approximation. We note that exact preimage computation is intractable for high dimensions, although a variant of this problem known as backward reachability has been studied in control [28] through exact computation or (guaranteed) over-approximation [24, 8]. The method of [16] is limited to preimage over-approximation and scales less well than PREMAP, and consequently our method. Compared to [3, 4], our divide-and-conquer approach focuses on minimizing the difference between the under-approximation and the preimage rather than maximizing a function value.
Methods of defending against patch attacks [31, 32] include active defences, such as small receptive fields or masking out all adversarial pixels from the input image, and passive defences, for example certifiable training using bound propagation. The most common of these, small receptive fields, extract features and predictions from small regions of the image, to avoid spatially localized attacks. The prediction involves robust aggregation such as majority voting. The smaller the receptive fields the better robustness but the worse clean performance, and [33] offers a computationally efficient trade-off.
Compared to active defences, which extend the model, our approach computes guarantees for the underlying neural network models. Certifiable training [19] involves estimating and propagating the bounds of neuron activations in each layer to bound the attacker’s influence on final predictions and is computationally expensive; in contrast, our method approximates the set of inputs for which the prediction is guaranteed for a given trained network.

### v2

@@ Abstract
The growing reliance on artificial intelligence in safety- and security-critical applications demands effective neural network certification. A challenging real-world use case is “patch attacks”, where adversarial patches or lighting conditions obscure parts of images, for example, traffic signs. A significant step towards certification against patch attacks was recently achieved using PREMAP, which uses under- and over-approximations of the preimage, the set of inputs that lead to a specified output, for the certification. While the PREMAP approach is versatile, it is currently limited to fully-connected neural networks of moderate dimensionality. In order to tackle broader real-world use cases, we present novel algorithmic extensions to PREMAP involving tighter bounds, adaptive Monte Carlo sampling, and improved branching heuristics. Firstly, we demonstrate that these efficiency improvements significantly outperform the original PREMAP and enable scaling to convolutional neural networks that were previously intractable. Secondly, we showcase the potential of preimage approximation methodology for analysing and certifying reliability and robustness on a range of use cases from computer vision and control. Our software is available in open source.
@@ 1 Introduction
The widespread deployment of AI technologies places significant demands on their trustworthiness and reliability, which are of particular concern in high-stakes applications, as recognized in the EU AI Act [1]. Much effort has been devoted to ensuring the safety and security of AI applications involving neural networks (NNs) due to their instability to adversarial perturbations [2]. A particularly challenging real-world example are “patch attacks” on images, such as traffic signs. In Figure 1, we show two examples of physical patch attacks: real graffiti on a traffic sign [3] (left) and a patch caused by a sunbeam [4] (middle).
@@
The examples above highlight the importance of studying the robustness of image recognition models to adversarial patch attacks for applications such as safe autonomous driving or biometric security software. To this end, a wide range of methodologies have been developed for neural networks to enhance their robustness, either via (certified) adversarial training [5, 6] or designing certifiable [7] or empirical [8] defences. An alternative method is to certify robustness by computing deterministic [9, 10] or probabilistic [11, 12, 13] guarantees on the network outputs.
Similarly to formal verification, certification is more powerful than adversarial training or empirical defences because it can guarantee that the network output is correct for all inputs in a given region. Using the patch attack in Figure 1 (right) as an illustration, certification aims to guarantee the prediction for all possible (RGB) colour combinations of the pixels in the black square.
@@
Many methods for adversarial robustness certification work by bounding the worst-case output of the neural network for a given input domain. An alternative approach, which gives more information than just the worst-case outputs, focuses on bounding the preimage of a given output specification (typically a polyhedron), either through over-approximating the preimage [14, 15, 16, 17] or under-approximating it [15, 16, 17]. Compared to over-approximation, under-approximation offers stronger compliance guarantees, since all inputs in the computed approximation satisfy the guarantee. Under-approximations also provide quantitative verification, i.e., estimating the proportion of inputs that meet the specification.
Several related methods [14, 17] rely on dividing the input space to obtain tighter approximations, which makes higher-dimensional inputs, such as images, intractable. Furthermore, non-linear activations are the cause of looser bounds [18] and convolutional neural networks (CNN) have significantly more non-linear activations compared to fully connected (FC) networks: even a small image has tens of thousands of non-linear activations in the first layer alone.
In this paper, we aim to broaden the applicability of the state-of-the-art preimage approximation algorithm PREMAP [15, 16] to a wider range of use cases. To this end, we improve the computational performance of PREMAP by significantly reducing the size of the computation tree through calculating tighter bounds, carefully selecting which neuron to branch on, and avoiding heavier optimizations when possible. These improvements enable preimage approximations of convolutional neural networks, which could not be handled previously. We demonstrate through empirical comparison that our method consistently outperforms the original PREMAP [16] on both under- and over-approximations, achieving running time improvements of at least an order of magnitude on reinforcement learning control tasks, while on patch attack certification the time-limited completion rates increase threefold for non-trivial cases.
We also showcase our method on a range of case studies: (i) certification for patch attacks on convolutional networks, which could not be handled previously, applied on GTSRB [19] (traffic signs), SVHN [20] (house numbers) and CIFAR [21] (standard image benchmark); (ii) out-of-distribution (OOD) input detection, where we show how to utilize both under- and over-approximations to certify whether inputs are outliers; and (iii) explainability (XAI), where we demonstrate how PREMAP can complement existing explainability techniques in analysing and debugging neural network models by highlighting important regions in the input. Our results demonstrate the potential of preimage approximation methodology for exploitation in reliability and robustness analysis and certification.
2 Related works
Robustness certification methods focus on computing deterministic [9, 10] or probabilistic/statistical [11, 12] guarantees on the output of a neural network. We can classify the methods into complete methods, such as constraint solving [10], or sound but incomplete methods, e.g., convex relaxation [22], which can be strengthened to completeness by employing branch-and-bound procedures [23, 24].
Building on [15, 16], our work employs convex (linear) relaxation and Monte Carlo sampling in a branch-and-refine framework for approximating the preimage of the neural network with exact (not statistical) quantitative guarantees. We leverage LiRPA [18] as implemented in CROWN tools [25, 26, 27] to compute symbolic relaxation bounds, adapting to preimage approximation.
We note that exact preimage computation is intractable for high dimensions, although a variant of this problem known as backward reachability has been studied in control [28] through exact computation or (guaranteed) over-approximation [29, 30]. Compared to [23, 24], our branch-and-refine approach focuses on minimizing the volume difference between the approximation and the preimage rather than maximizing a function value. The methods of [14, 17] also perform preimage approximations, but the former only over-approximations. However, they both [14, 17] use input-space splitting, which scales poorly beyond low-dimensional data. Methods defending against patch attacks [8, 31] include active defences, such as small receptive fields or masking out all adversarial pixels from the input image, and passive defences, for example, certifiable training using bound propagation. The most common of these, small receptive fields, extract features and predictions from small regions of the image, to avoid spatially localized attacks. The prediction involves robust aggregation such as majority voting. The smaller the receptive fields the better robustness but the worse clean performance, where [7] offers a computationally efficient trade-off. Compared to active defences, which extend the model, our approach computes guarantees for existing neural network models.
Certifiable training [6] involves estimating and propagating the bounds of activations in each layer to bound the attacker’s influence on final predictions and is computationally expensive. Randomized smoothing [13] consists of repeated predictions with random noise added to the input during inference or training [32], which yields probabilistic certification of robustness. In contrast, our method provides quantitative certification of robustness by approximating the set of inputs for which the prediction is guaranteed for a given network.
3 Preliminaries
@@
In this section, we define preliminary concepts and provide an overview of the PREiMage APproximation (PREMAP) method [15, 16], focusing on key aspects needed for our algorithmic improvements; for more detail please refer to [16]
3.1 Neural networks
@@
We use $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ to denote a feed-forward neural network, where $f(\mathbf{x})=\mathbf{y}$. For layer $l$, $\mathbf{W}^{l}$ denotes the weight matrix, $\mathbf{b}^{l}$ the bias, $\mathbf{z}^{l}$ the pre-activation neurons, and $h^{l}(\mathbf{z}^{l})$ the activation function, such that $\mathbf{z}^{l}=\mathbf{W}^{l}h^{l-1}(\mathbf{z}^{l-1})+\mathbf{b}^{l}$. We use under- and over-lines to denote the lower and upper bounds, e.g., $\underline{\mathbf{x}}_{i}\leq\mathbf{x}_{i}\leq\overline{\mathbf{x}}_{i}$. In this paper we focus on ReLU neural networks, where $h^{l}(\mathbf{z}^{l})=\text{ReLU}(\mathbf{z}^{l})=\max(\mathbf{z}^{l},0)$ is applied element-wise, but our method can also be generalized to other activation functions that can be bounded by linear functions [25, 33].
3.2 Linear relaxation
@@
For the neural network $f$, linear relaxation is used to compute linear lower and upper bounds for the output:
@@ 3.4 PREMAP
The PREMAP algorithm [16, 15], implements both under- or over-approximation of the preimage of a given output specification $\mathbb{O}$ restricted to the input domain $\mathbb{I}$, in the form of a disjoint union of polytopes. Algorithm 1 shows an overview of the PREMAP method with our modifications highlighted with green shading (discussed further in Section 4). For brevity, some equations are only shown for under-approximations, but work equivalently for over-approximations. PREMAP relies on linear relaxation with LiRPA to approximate the preimage (see line 3), which is further refined through splitting the problem into disjoint subregions (on 8 and 10), that can be approximated independently in parallel (see 12 and 13) with a branch-and-refine approach. The algorithm can be terminated anytime, but stops upon reaching a preset threshold for the preimage approximation.
3.5 Branching and refinement
@@
Since the initial bound given by LiRPA might be very loose (Algorithm 1, line 3), PREMAP iteratively partitions a subdomain $\mathbb{I}_{\text{sub}}\subseteq\mathbb{I}$ into two smaller subdomains $\mathbb{I}_{\text{sub}}^{-}$ and $\mathbb{I}_{\text{sub}}^{+}$ (5 to 14). We select which subdomain to split by the largest difference in estimated volume between the exact preimage and the approximation in the subdomain (line 6). The split can be either on the input, which partitions the input into two subdomains along some feature, or based on the pre-activation value of an intermediate unstable neuron [16], called ReLU splitting. As this paper deals with larger input domains, for which input splitting performs less well, we will focus on the latter (line 8).
@@
A key advantage of ReLU splitting is that it allows us to replace unstable neuron bounds with precise bounds. For an unstable ReLU neuron $h^{l}(\mathbf{z}^{l}_{j})=\max(0,\mathbf{z}^{l}_{j})$, we use linear relaxation to bound the post-activation value as in Equation 1. When a split is introduced, the neuron becomes stable in each subdomain, and the exact linear functions $h^{l}(\mathbf{z}^{l}_{j})=\mathbf{z}^{l}_{j}$ and $h^{l}(\mathbf{z}^{l}_{j})=0$ can be used in place of its linear relaxation. This typically tightens the approximation on each subdomain as the linear relaxation errors for this unstable neuron are removed and substituted with the exact symbolic function.
@@
The bounds, $\underline{\mathbf{A}}\mathbf{x}+\underline{\mathbf{b}}$, for the new under-approximations (or $\overline{\mathbf{A}}\mathbf{x}+\overline{\mathbf{b}}$ for over-approximations) are optimized via projected gradient descent to maximize the approximation (Algorithm 2, line 3) and the process of refining the subdomains continues until an appropriate termination criterion is met (Algorithm 1, line 7). The procedure can also be stopped anytime to produce a valid (but looser) approximation. The union of the preimage approximations $\mathbf{P}$ for the disjoint subdomains $\mathbb{I}_{\text{sub}}$ constitutes the preimage approximation of the output property.
@@
2:  $\mathbf{X}\leftarrow\textsc{Sample}(\mathbb{I},5n)$ $\rangle\rangle$ Sample extra when it is easy
@@
4:  $\mathbb{D}\leftarrow\{(\mathbf{P},\mathbb{I},\mathbf{X})\}$ $\rangle\rangle$ Approximation, domain, samples
5:  while $\textsc{EstimateApproxRatio}(\mathbb{D},f_{\mathbb{O}})<v_{\text{th}}$
@@
7:   $\mathbf{X}\leftarrow\textsc{Sample}(\mathbb{I}_{\text{sub}},n,\mathbf{X})$ $\rangle\rangle$ See Section 4.2
8:   $l,i\leftarrow\textsc{SelectNeuron}(f_{\mathbb{O}},\mathbb{I}_{\text{sub}},\mathbf{X})$ $\rangle\rangle$ Section 4.3
9:   $\mathbb{I}_{\text{sub}}\leftarrow\textsc{CheckShortcut}(\mathbb{I}_{\text{sub}},\mathbf{X})$ $\rangle\rangle$ See Section 4.4
10:   $\mathbb{I}_{\text{sub}}^{-},\mathbf{X}^{-},\mathbb{I}_{\text{sub}}^{+},\mathbf{X}^{+}\leftarrow\textsc{SplitNeuron}(\mathbb{I}_{\text{sub}},\mathbf{X},l,i)$
11:   $\textsc{PropagateBounds}(\mathbb{I}_{\text{sub}}^{-},\mathbb{I}_{\text{sub}}^{+},l,i)$ $\rangle\rangle$ See Section 4.1
12:   $\mathbf{P}^{-}\leftarrow\textsc{PreimageApprox}(f_{\mathbb{O}},\mathbb{I}_{\text{sub}}^{-},\mathbf{X}^{-})$
13:   $\mathbf{P}^{+}\leftarrow\textsc{PreimageApprox}(f_{\mathbb{O}},\mathbb{I}_{\text{sub}}^{+},\mathbf{X}^{+})$
14:   $\mathbb{D}\leftarrow\mathbb{D}\cup\{(\mathbf{P}^{-},\mathbb{I}_{\text{sub}}^{-},\mathbf{X}^{-}),~(\mathbf{P}^{+},\mathbb{I}_{\text{sub}}^{+},\mathbf{X}^{+})\}$
15:  return $\{\mathbf{P}\mid(\mathbf{P},\cdot,\cdot)\in\mathbb{D}\}$
Algorithm 1: Premap is an algorithm for finding preimage under approximations. The inputs are the neural network $f$, the input domain $\mathbb{I}$, the output specifications $\mathbb{O}$, the early stopping threshold $v_{\text{th}}$, and the number of samples $n$. Some functions are detailed in Algorithm 2. The lines with improvements compared to [16] are highlighted with green shading and are discussed in Sections 4.2, 4.1 and 4.3.
@@
3:  $\boldsymbol{\alpha}\leftarrow\operatorname*{arg\,max}\limits_{\boldsymbol{\alpha}}\sum\limits_{\mathbf{x}\in\mathbf{X}}\sigma(-\log(\sum\exp({-\underline{\mathbf{A}}^{(\boldsymbol{\alpha})}\mathbf{x}-\underline{\mathbf{b}}^{(\boldsymbol{\alpha})}})))$
@@
5:  return $\mathbf{P}$ $\rangle\rangle$ Preimage under approximation
6:function EstimateApproxRatio($\mathbb{D},f_{\mathbb{O}}$)
@@
9:  return $v_{\mathbf{P}}/v_{\mathbb{O}}$ $\rangle\rangle$ Approximated fraction of the preimage
@@
12:  return $v_{\mathbb{O}\setminus\mathbf{P}}$ $\rangle\rangle$ Volume of preimage not approximated
13:function SplitNeuron($\mathbb{I}_{\text{sub}},\mathbf{X},l,i$)
14:  $\mathbb{I}_{\text{sub}}^{-},\mathbf{X}^{-}\leftarrow\{\mathbf{x}\mid\mathbf{x}\in\mathbb{I}_{\text{sub}}\wedge\mathbf{z}^{l}_{i}<0\},\{\mathbf{x}\mid\mathbf{x}\in\mathbf{X}\wedge\mathbf{z}^{l}_{i}<0\}$
15:  $\mathbb{I}_{\text{sub}}^{+},\mathbf{X}^{+}\leftarrow\{\mathbf{x}\mid\mathbf{x}\in\mathbb{I}_{\text{sub}}\wedge\mathbf{z}^{l}_{i}\geq 0\},\{\mathbf{x}\mid\mathbf{x}\in\mathbf{X}\wedge\mathbf{z}^{l}_{i}\geq 0\}$
16:  return $\mathbb{I}_{\text{sub}}^{-},\mathbf{X}^{-},\mathbb{I}_{\text{sub}}^{+},\mathbf{X}^{+}$ $\rangle\rangle$ Split domain and samples
Algorithm 2: Additional functions for Premap, see Algorithm 1 for details. The Lirpa[18] function produces linear bounds with optimizable parameters $\boldsymbol{\alpha}$. And $\mathbf{z}^{l}_{i}$ is the activation of neuron $i$ on layer $l$.